OpenStudy (anonymous):

6 – 4(3 – 5)^2 + 30 ÷ 5 6 – 4(–2)^2 + 30 ÷ 5 6 – 4(4) + 30 ÷ 5 6 – 16 + 30 ÷ 5 −10 + 30 ÷ 5 20 ÷ 5 Matt's answer 4

OpenStudy (camerondoherty):

Matt's Answer is wrong its supposed to be -4

OpenStudy (igreen):

The answer is supposed to be -4.

OpenStudy (igreen):

Lol.

OpenStudy (anonymous):

thanks

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. \(\bbox[5pt, #000000 ,border:5px solid #00FF00]{\bf\huge \color{#00FF00}\sim \! \! \! \! \color{#00FF00}i\it\Huge\color{#00FF00}G\color{#00FF00}r\color{#00FF00}e\color{#00FF00}e\color{#00FF00}n\bf\color{#00FF00}ツ}\)

OpenStudy (igreen):

Though I think you have to go to the step in which he was wrong..

OpenStudy (camerondoherty):

Step 1 : Raise -2 to the 2nd power A Negative number raised to an even power is positive For example let's look at (-7)6 , where (-7) , a negative number, is raised to 6 , an even exponent : (-7)6 can be written as (-7)•(-7)•(-7)•(-7)•(-7)•(-7) Now, using the rule that says minus times minus is plus, (-7)6 can be written as (49)•(49)•(49) which in turn can be written as (7)•(7)•(7)•(7)•(7)•(7) or 76 which is positive. We proved that (-7)6 is equal to (7)6 which is a positive number Using the same questions as above, replacing (-7) by any negative number, and replacing the exponent 6 by any even exponent, we proved which had to be proved Equation at the end of step 1 : \[ (6 - (4 \times 22)) + 6 \] Step 2 : Multiply 4 by 22 Multiplying exponents : 22 multiplied by 22 = \[ 2(2 + 2) = 24 \] Equation at the end of step 2 : \[ (6 - 24) + 6 \] Step 3 : Simplify -10 + 6 Final result : -4

OpenStudy (igreen):

It would be step 3.

OpenStudy (igreen):

Err..step 4. 4 * 4 = 16, NOT -16.

geerky42 (geerky42):

-4(4) = -16...